Automobile batteries are filled with an aqueous solution of sulfuric acid. What is the mass of the acid (in grams) in 900. mL of the battery acid solution if the density of the solution is 1.285 and the solution is 38.08% sulfuric acid by mass?
1
Expert's answer
2016-10-29T11:04:08-0400
ρ=m/V
m= ρ · V

m (H2SO4 38.08%)=1.285 · 900 = 1156.5 g
m (H2SO4 100%)=1156.5 · (38.08 / 100) = 440.39 g